OpenStudy (anonymous):

Solve 8x-4(-5x-2)=12x

OpenStudy (saifoo.khan):

8x + 20x +8 = 12x 28x - 12x + 8 = 0 16x + 8 = 0 x = -8/16 x = -1/2